OpenStudy (anonymous):

sqrt3sinx +cosx=sqrt3

OpenStudy (kc_kennylau):

Hint: sin x = 1 and cos x = 0

OpenStudy (raden):

for asin(x) + bcos(x) = c, the left side can be converted to k cos(x - θ) with k = sqrt(a^2+b^2) and to get θ use tanθ = a/b, then solve for θ so that the equation would be kcos(x-θ) = c then find the solutions for x

OpenStudy (anonymous):

Divide both side by V3 (sqr.3) sin x + cos x/V3 = 1. Replace in the equation 1/V3 = tan Pi/6 = = (sin Pi/6)/(cos Pi/6). sin x*cos Pi/6 + sin Pi/6*cos x = 1 = sin Pi/2 Transform the left side by using the trig identity: sin (a + b) = sin a*cos b + sin b*cos a sin (x + Pi/6) = sin Pi/2 x + Pi/6 = Pi/2 -> x = Pi/2 - Pi/6 = 2Pi/6 = Pi/3